# Retrieval-Augmented Generation: Is Dense Passage Retrieval Retrieving?

### v1

@@ 8 Limitations
This paper presents a detailed analysis of the DPR formula, specifically focusing on the original DPR training formula utilizing a BERT backbone. We anticipate that our findings will exhibit a degree of generalizability across various DPR implementations, given the underlying commonalities of the core training approach. It is important to recognize that modifications—such as improving hard negatives, different data augmentation techniques, different transformer-based backbones, or leveraging multiple views/vectors from models—while serving to refine and enhance the DPR framework, build upon and amplify the mechanisms of the DPR method. These enhancements, though significant in optimizing performance, are expected not to fundamentally change this analysis. However, it is still a limitation of this paper that we did not repeat our analysis on more DPR-based methods and datasets.
@@ A.2 Model Activations at different thresholds
Figures 6, 7, 8, 9, and 10 illustrate neuron activation patterns across varying activation thresholds set at $0.005*\max(Attr)$, $0.01*\max(Attr)$, $0.05*\max(Attr)$, $0.2*\max(Attr)$, and $0.3*\max(Attr)$, respectively. As the threshold increases from 0.005 to 0.3, the visualization narrows down to neurons with stronger activations. This observation reinforces the findings discussed in Section 4: pre-trained BERT shows a trend of fewer but more consistently activated neurons, in contrast to DPR-trained BERT, which exhibits a broader array of neurons activated less frequently.

### v2

@@ 7 Limitations
This paper presents a detailed analysis of the DPR formula, specifically focusing on the original DPR training formula utilizing a BERT backbone. We anticipate that our findings will exhibit a degree of generalizability across various DPR implementations, given the underlying commonalities of the core training approach. It is important to recognize that modifications—such as improving hard negatives, different data augmentation techniques, different transformer-based backbones, or leveraging multiple views/vectors from models—while serving to refine and enhance the DPR framework, build upon and amplify the mechanisms of the DPR method. These enhancements, though significant in optimizing performance, are expected not to fundamentally change this analysis. However, it is still a limitation of this paper that we did not repeat our analysis on more DPR-based methods and datasets. Reichman and Heck (2024)
@@ A.2 Model Activations at different thresholds
Figures 3, 4, 5, 6, and 7 illustrate neuron activation patterns across varying activation thresholds set at $0.005*\max(Attr)$, $0.01*\max(Attr)$, $0.05*\max(Attr)$, $0.2*\max(Attr)$, and $0.3*\max(Attr)$, respectively. As the threshold increases from 0.005 to 0.3, the visualization narrows down to neurons with stronger activations. This observation reinforces the findings discussed in Section 3: pre-trained BERT shows a trend of fewer but more consistently activated neurons, in contrast to DPR-trained BERT, which exhibits a broader array of neurons activated less frequently.